If the minute hand of a big clock is 1.05 m long, find the rate at which its tip is moving in cm per minute. 

Grade:10

1 Answers

Agastya Varma
6 years ago
Length of minute hand = 1.05cm =105cm
To find the rate at which it is moving we have to find distance(arc length) covered by it in 1 min
Circumference =2(pi)r = 2*22/7*105 = 660cm
So, the hand travels 660cm in one hour or 60 min
So, in one minute it would travel (660/60) = 11cm/min
Hope it helped you